### v1

@@ 5 Numerical results
In this section, we extensively compare several MPECopt variants to some of the most successful regularization and penalty-based methods on two distinct nonlinear MPEC benchmarks.
@@ 5.1 Implementation
[…]
$x_{1}-g(x)=0,$
$x_{2}-h(x)=0,$
$0\leq x_{1}\perp x_{2}\geq 0$
[…]
@@
We use two different ways to solve the LPECs. First, we use the MILP formulation in (21) and solve it with the commercial solver Gurobi Gurobi or the open-source solver HiGHS Huangfu2018 . In our experiments, we used Matlab R2024a, where the HiGHS solver can be directly called from Matlab via the intlinprog function. For the MILPs, we set the maximum number of nodes in the branch and bound algorithm to 500 and the maximum computation time to 5 minutes. The tolerance for the LP solves in both MILP solvers is set to $10^{-9}$ (default is $10^{-6}$). Second, as LPECs are a special case of MPECs, we solve them via the homotopy regularization and penalty-based methods described below. The homotopy terminates if the solver reports local infeasibility or if a complementarity residual of $\|\mathrm{diag}(x_{1})x_{2}\|_{\infty}$ is below $10^{-9}$. The subproblems are solved with IPOPT.
For comparisons, we also implement the regularization and penalty-based methods from Section 2.2. Thereby, an adaptation of Reg($\tau$) and Pen($\tau$) to the form of  (38) are solved in a homotopy loop with a decreasing parameter $\tau$. We initialize $\tau^{0}=1$ and update it iteratively using the rule $\tau^{k}=0.1\tau^{k-1}$. The loop terminates if either: the NLP solution is feasible and the complementarity tolerance $h_{\perp}(x)$ is below $10^{-9}$; an NLP is locally infeasible; or a maximum number of $15$ iterations is reached. A homotopy loop is declared successful if it finds a stationary point. Recall that within the homotopy approach, unless the identified point is S-stationary, it remains unclear whether it is also B-stationary.
Our implementation of MPECopt is open-source and available at https://github.com/nosnoc/mpecopt. The repository contains several tutorial examples and the two benchmarks below.
@@ 5.2 Results on MacMPEC benchmark
We explore several algorithmic options of MPECopt on the MacMPEC benchmark set Leyffer2000 . The MacMPEC problem set is available in the form of a package of AMPL .mod and .dat files from wiki.mcs.anl.gov/leyffer/index.php/MacMPEC. We consider in total 191 problems. They are translated into the CasADi symbolic framework and are available in https://github.com/nosnoc/mpecopt/benchmarks/macmpec.
5.2.1 Comparison of different Phase I algorithms
In our first experiment, we compare several Phase I algorithms discussed in Section 3.3. In particular, we examine the following five versions:
  1. 1.
    Algorithm 2 with Reg($\tau$) and LPEC solves, $\rho_{I}=10^{-1}$eg-LPEC.
  2. 2.
    Algorithm 2 with Reg($\tau$) and simple projection in (24) – Reg-Simple.
  3. 3.
    Algorithm 2 with $\ell_{1}$ reformulation of Pen($\tau$) and LPEC solves, $\rho_{I}=10^{-1}$en-$\ell_{1}$-LPEC.
  4. 4.
    The $\ell_{1}$ reformulation of (25) via Algorithm 3, $\rho^{0}=10^{0}$easibility-$\ell_{1}$.
  5. 5.
    The $\ell_{\infty}$ reformulation of (25) via Algorithm 3, $\rho^{0}=10^{0}$$\ell_{\infty}$.
Algorithm 2 with an $\ell_{\infty}$, instead of $\ell_{1}$ reformulation, achieves similar results as the $\ell_{1}$-approach, while being only slightly slower. We omit it from the plots for the sake of better readability. For all versions, we tried several initial trust region radii with smaller and larger values than in the list above. For the sake of brevity, we picked for each variant the best value and report only these results here. The initial trust region did not have much influence on the regularization-based methods within Algorithm 2, and the feasibility formulation (25) had slightly higher success rates with a larger value of the initial radius. In this experiment, Gurobi is used for solving LPECs by rewriting them into MILPs (21).
Results of the experiments are given in the Dolan-Moré absolute and relative performance profiles Dolan2002 . Phase I is declared as successful if it finds a feasible point $x^{0}\in\Omega$. The results for the different Phase I algorithms are given in Fig. 6.
@@ 5.2.1 Comparison of different Phase I algorithms
Figure 6: Evaluating different Phase I algorithms in MPECopt on the MacMPEC test set.
Overall, Reg-LPEC and Reg-Simple are the most successful approaches with a success rate of 95.29%, but they are not always the fastest. The simple projection, Reg-Simple, is mostly very effective for easy MPECs where there are no biactive constraints, as it quickly finds a feasible BNLP and does not require LPEC solves. However, it requires significantly more NLP solves on some examples, independent of the number of biactive constraint at the solution, and requires many NLP solves until $\tau$ becomes very small, which is needed to obtain a valid active set estimate via (24). In the worst case, on several problem instances such as pack-comp1-8, pack-comp1c-8, pack-comp1p-8, it requires 20 NLP solver calls, whereas solves Reg-LPEC require at most 8 NLP solves to compute a B-stationary point. In such cases, Reg-LPEC is faster, as it can be seen in Fig. 6. Pen-$\ell_{1}$-LPEC is often the fastest approach but has a lower success rate with 93.19%. The reason is that penalty strategies are in general faster than the regularization-based approach, but have a lower success rate in finding any stationary point of an MPEC. This is consistent with the overall results for these two strategies, reported in Fig. 10. However, even to Pen-$\ell_{1}$-LPEC is relatively the fastest, in Fig. 6a it can be seen that in absolute terms it is not significantly faster.
The feasibility-based methods $\ell_{\infty}$ and Feasibility-$\ell_{1}$ are in this benchmark less competitive than Algorithm 2, with success rates of 79.05% and 75.39%, respectively. In all other cases when they fail, a B-stationary point of  (25) with a nonzero residual is found.
Another advantage of Algorithm 2 over the feasibility approach is that they already take the objective of (1) into account. Therefore, the first feasible BNLP found Algorithm 2 and its solution $x^{0}$ may already yield a B-stationary point of the MPEC. In such cases, the first LPEC solve in Algorithm 3 only needs to verify B-stationarity. For Reg-LPEC this happens in 56.02%, for Reg-Simple in 74.34% and $\ell_{\infty}$-LPEC in 16.23% of the overall successful cases. In other words, when the regularization-based method is successful in finding B-stationary points, MPECopt inherits its efficacy, and Phase II immediately certifies the success. Moreover, an LPEC can often identify a feasible BNLP even for larger values of $\tau$. This implies a smaller number of total NLP solves in MPECopt compared to a regularization-based method, which we discuss in more detail in Section 5.2.3.
In conclusion, Reg-LPEC is the most successful Phase I approach and our default option for all other experiments below. As it can be seen in the absolute timing in Fig. 6a it is only slightly slower than Reg-Simple on easier MPECs, but manages to find reliably a feasible point in most of the cases. Recall that, although Reg-Simple does not require solving LPECs in Phase I, it still needs to solve LPECs in Phase II to verify B-stationarity or to determine a new BNLP.
(a) Absolute timings.
(b) Relative timings.
Figure 7: Evaluating different LPEC algorithms in MPECopt on the MacMPEC test set.
5.2.2 Comparison of different LPEC solvers
In our next experiment, we compare different LPEC solution strategies. In all experiments, we use the regularization-based Phase I Reg-LPEC, but we use different solvers for solving the LPECs in both phases. For the comparison, we consider four distinct variants where the LPEC is solved via
  1. 1.
    the MILP reformulation  (21) with Gurobi, and $\rho_{I}=10^{-1}$ in Phase I – Gurobi-MILP.
  2. 2.
    the MILP reformulation  (21) with HiGHS, and $\rho_{I}=10^{-3}$ in Phase I – HiGHS-MILP.
  3. 3.
    Generic MPEC with regularization-based reformulation Reg($\tau$)Reg-MPEC.
  4. 4.
    Generic MPEC with $\ell_{1}$-penalty reformulation Pen($\tau$$\ell_{\infty}$-MPEC.
The HiGHS solver performs significantly better if the trust region in the LPEC is smaller. For this reason, we set in Phase I for this solver $\rho_{I}=10^{-3}$. Gurobi is less sensitive to these changes and has similar performance for larger and smaller values of $\rho_{I}$. We also conducted experiments using the $\ell_{\infty}$-penalty method, which produced results similar to those of the $\ell_{1}$-penalty method. To improve readability, we have omitted the $\ell_{\infty}$-penalty results from the plots blow.
We solve all problems in MacMPEC with the four different LPEC solution strategies. Thereby, we compare the total cumulative computation times spent for solving all LPECs in each MPEC instance of MacMPEC. In 34 of the 191 problems, after the first NLP is solved, either the MPEC is already solved or found to be locally infeasible. We exclude these from the plots, as the effective LPEC wall time in these 34 problems is zero. The remaining 157 problems are successfully solved by all approaches, except for HiGHS-MILP, which does not solve LPECs appearing in the problems siouxfls and siouxfls1.
Fig. 7 shows the absolute and relative performance profile with the total cumulative LPEC solution times. It can be seen that MILP strategies significantly outperform the regularization and penalty methods in terms of computation time.
This large difference has several causes. The MILP methods solve the LPEC to global optimality (which is sufficient for verifying B-stationary), whereas the other can compute only a stationary point (which may not even be a local minimum or B-stationary point of the LPEC). Consequently, LPEC solved by regularization-based methods tend to be more often locally infeasible in Phase I or need to solve more LPECs until a B-stationary point is verified in Phase II. To have a more fair comparison, in addition to the cumulative times, in Fig. 8 we show the maximum LPEC computation time in every of the 157 MPEC instances. Even there, we see that the MILP-based methods are up to two orders of magnitude faster than regularization-based methods for LPECs on the MacMPEC test set. This empirically confirms the arguments made in Section 3.2.2 on efficient solving LPECs as MILPs. Moreover, we observed often that, when the solution is S-stationary, the MILP is solved after one LP solve, which complies with the result of Lemma 3.3.
Figure 8: Maximal solution times for different LPEC algorithms in MPECopt on the MacMPEC test set.
Finally, we compare the cumulative NLP and LPEC computation times in every MPECopt solver call on every MacMPEC problem. The results are reported in Fig. 9, where the diagonal line means equal time spent on NLP and LPEC solving. We see that for the MILP methods, most dots are far below the diagonal, which means that a smaller fraction of the total computation time is spent on solving the LPECs, and the expensive part is solving the BNLPs. On the other hand, in the regularization-based methods, the dots cluster around and above the diagonal, meaning that the total computational burden is almost equally spread on solving NLPs and LPECs. In particular, the cumulative NLP time is smaller than the cumulative LPEC time in 0% of cases for Gurobi-MILP, 20% for HiGHS-MILP, 87.26 % for Reg-MPEC, and 85.99 % for $\ell_{\infty}$-MPEC. The MILP solvers, especially HiGHS profit from smaller initial true region radii in Phase I. However, the smaller $\rho_{I}$ the more NLPs in Phase I must be solved until a feasible LPEC is found, which shifts the computational load more to the NLP side.
To conclude, our experiments on MacMPEC confirm that solving LPECs via MILPs can be done very efficiently in practice, both with commercial (Gurobi) and open-source (HiGHS) solvers. Remarkably, in this benchmark, the MILP approaches are often orders of magnitude faster than the integer-free regularization-based methods for solving LPECs.
5.2.3 Comparison of MPECopt to regularization-based methods
(a) Absolute timings.
(b) Relative timings.
@@ 5.2.3 Comparison of MPECopt to regularization-based methods
Figure 10: Evaluating different MPEC solution methods on the MacMPEC test set in terms of finding a stationary point.
Next, we compare MPECopt to the regularization and penalty-based solution methods from Section 2.2. For Phase I of MPECopt, we use again Reg-LPEC, which had the highest success rate in experiments from Section 5.2.1. In the following experiment, we compare in total five algorithms:
@@
    MPECopt, with $\ell_{1}$-penalty-based Phase I, and Gurobi as LPEC solver – MPECopt-$\ell_{1}$-Gurobi.
@@
    MPECopt, with regularization-based Phase I, and HiGHS as LPEC solver – MPECopt-HiGHS.
@@
    Regularization-based method, that solve Reg($\tau$) in a homotopy loop– Reg.
@@
    $\ell_{1}$-penalty-based method, that solve an $\ell_{1}$ reformulation of Pen($\tau$) in a homotopy loop – Pen-$\ell_{1}$.
@@
We also tried the $\ell_{\infty}$ penalty method, which similar computation time as $\ell_{1}$, but a slightly lower success rate. Therefore, for better readability, we omit it from the plots. Moreover, MPECopt with the simple projection strategy, i.e., Reg-Simple from Section 5.2.1 performs similar to MPECopt-Reg-Gurobi, and we omit it as well for better readability of the plots.
@@
We first compare the success rate and computations point for finding a stationary point. Note that, if MPECopt is successful, it always returns a B-stationary point, whereas for the other methods, we can only know that it is a B-stationary point if it is also S-stationary. Fig. 10 reports the absolute and relative performance profiles for this experiment.
We can see that MPECopt-Reg-Gurobi achieves in total the highest success rate of 94.24%. It is faster and more successful than its counterpart Reg (92.67%). The $\ell_{1}$ penalty method is faster than MPECopt-Reg-Gurobi, but with a smaller success rate of 92.14%. On the other hand, MPECopt-$\ell_{1}$-Gurobi is only slightly slower than Pen-$\ell_{1}$ and achieves the same success rate. MPECopt-Reg-HiGHS is slightly slower than MPECopt-Reg-Gurobi, and has a lower success rate of 92.67%, because HiGHS fails to solve LPECs in Phase II in the problems incid-set-32, incid-set2-32 and incid-setc-32. The difference in computation times is consistent with Fig. 9.
Figure 11: Number of NLPs (top plot) and NLPs (bottom plots) solved in MacMPEC on all problem instances.
The MPECopt method inheres the efficacy of the standard methods used in Phase I, but can often be even faster as it requires fewer NLP solves. The reason is that already for a moderate value of $\tau$, the LPEC can identify a feasible BNLP in Phase I. Figure 11 shows the number of NLPs and LPECs solved in every problem instance. The plot does not include the NLP and LPEC solved used to determine the type of stationary points for Reg and Pen-$\ell_{1}$ in the next section. We observe that MPECopt-Reg-Gurobi and MPECopt-Reg-HiGHS consistently solve a smaller number of NLPs than their counterpart Reg. Similar holds for MPECopt-$\ell_{1}$-Gurobi and Pen-$\ell_{1}$. Moreover, MPECopt requires in most cases only a few LPEC solutions. In many cases, it is only two, one in Phase I to identify a feasible BNLP, and one in Phase II to verify B-stationarity.
All methods are quite fast on easy problems where there are no biactive constraints, and often a single NLP solve is sufficient to find a solution. Approximately more than two-thirds of solutions found by any of the algorithms have an empty biactive set.
5.2.4 Comparison of stationary points
Figure 12: Distribution of stationary points on MacMPEC for different solution methods. Failure counts the number of infeasible problems.
Now we look at the distribution of stationary point types found by the algorithms compared in the previous section. We also check are these points B-stationary. To determine the type of stationary point from Definition 2.3 we must solve the TNLP (9), which requires a correct identification of the active sets ${\mathcal{I}}_{+0},{\mathcal{I}}_{0+}$ and ${\mathcal{I}}_{00}$. MPECopt is an active set method where the complementarity constraints are usually satisfied exactly or close to machine precision, making it easy to find the active sets for the TNLP. On the other hand, Reg, and Pen-$\ell_{1}$ rarely have a complementarity residual $h_{\perp}(x)$ much smaller than the prescribed tolerance of $10^{-9}$, making an active set identification more difficult. We take an iterative procedure, and identify the active set via Eq. (19), with $\epsilon=10^{-6}$. To minimize the possibility of wrong conclusions, we implement several safeguards. We denote the solution computed via Reg or Pen-$\ell_{1}$ by $x^{*}$, and the corresponding TNLP solution by $x^{\mathrm{tnlp}}$. We declare the active set to be correctly identified if $\|x^{*}-x^{\mathrm{tnlp}}\|\leq 10^{-7}$. If this is not satisfied we reduce $\epsilon$ by a factor of 10 and repeat the procedure. Once $\|x^{*}-x^{\mathrm{tnlp}}\|\leq 10^{-7}$ is true or a maximum number of 5 attempts is exceeded, we compute the multipliers in the TNLP and classify the stationary point. To verify B-stationarity, we take several steps. First, we use the point $x^{\mathrm{tnlp}}$ and construct an ${\mathrm{LPEC}}(x^{\mathrm{tnlp}},\rho)$ with $\rho=10^{-3}$. If the objective of this LPEC is below $\texttt{tol}_{\mathrm{B}}=10^{-8}$, then we declare a $x^{*}$ to be B-stationary. If this is not the case, we take up to three more LPEC solve attempts while reducing $\rho$ by a factor of 10. The minimal value is $\rho=10^{-6}$, which is the minimal allowed trust region radius in MPECopt. This delivers a valid certificate in most cases. To have more robustness to wrong active set identification, if this does not verify B-stationarity, we repeat the same but now with $x^{*}$ as linearization point ${\mathrm{LPEC}}(x^{*},\rho)$. Finally, if this also fails, we look at the objective $f(x^{*})$ and the objective computed by MPECopt for the same problem, which is denoted by $f(\hat{x})$. If the relative difference satisfies $\frac{|f(x^{*})-f(\hat{x})|}{|f(\hat{x})|+10^{-16}}<10^{-3}$, then we declare $x^{*}$ to be B-stationary, if $\hat{x}$ is B-stationary. If the difference between the objective is larger, and the LPECs find a descent direction, the point $x^{*}$ is declared to be not B-stationary.
Figure 12 plots the distribution of the stationary points for all solvers except MPECopt-Reg-HiGHS, which is the same as MPECopt-Reg-Gurobi (except on the problems that were not solved). S stationary points dominate the overall picture, as two-thirds of problems have no biactive constraints. MPECopt always converges to a B-stationary point by construction. Using the above-described procedure for identifying stationary points, Reg and Pen-$\ell_{1}$ fail only in two cases to find a B-stationary point. This happens on the problems ralph2 (A-stationary) and scholtes4 (C-stationary). MPECopt computes in both cases a B-stationary point with the objective value of $10^{-12}$, whereas Reg computes objectives of $0.6325\cdot 10^{-4}$, which is too far away from the B-stationary point. Similar results are obtained with Pen-$\ell_{1}$. This can be improved by reducing significantly the complementarity tolerance in the homotopy loop, e.g. $10^{-14}$ instead of $10^{-9}$. However, this makes Reg and Pen-$\ell_{1}$ significantly slower, and fail more often on other problems.
@@ 5.2.4 Comparison of stationary points
In summary, we can say that all methods, sometimes with some tuning, can successfully compute B-stationary points on MacMPEC (and MPECopt directly certifies this). To get a broader perspective, we run another benchmark in the next section.

### v2

@@ 5 Numerical results
In this section, we extensively compare several MPECopt variants with regularization-based methods, as well as with solving the MINLP and NLP reformulations of MPECs on two distinct nonlinear MPEC benchmarks.
@@ 5.1 Implementation
where $c^{\mathrm{lb}},c^{\mathrm{ub}}\in{\mathbb{R}}^{n_{c}}$ are lower and upper bounds of $c(x)$ with $c^{\mathrm{ub}}_{i}\geq c^{\mathrm{lb}}_{i},\ \forall i\in\{1,\ldots,n_{c}\}$. Equality constraints are imposed by settings $c^{\mathrm{ub}}_{i}=c^{\mathrm{lb}}_{i}$. The vectors $x^{\mathrm{lb}},x^{\mathrm{ub}}\in{\mathbb{R}}^{n}$ are lower and upper bounds on $x$. If there is no lower or upper bound on some $x_{i}$ or $c_{i}(x)$, then we set the respective value on $\texttt{inf}$ or inf. The functions $g,h:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ are assumed twice continuously differentiable.
To solve the NLPs arising in MPECopt we use the IPOPT solver Waechter2006 . For better performance, we change some of the default options by setting the tolerance to $\texttt{tol}=10^{-12}$ (default is $10^{-8}$), $\texttt{acceptable\_tol}=10^{-9}$ (default is $10^{-6}$), and $\texttt{warm\_start\_init\_point}=\texttt{'yes'}$. Moreover, we change the barrier parameter strategy by setting: mu_strategy = ’adaptive’ and mu_oracle = ’quality-function’, cf. Nocedal2009 . In all experiments, we use MA27 Duff1982 from the HSL library HSL as the linear solver in IPOPT. The same settings are used for all methods and NLP solutions in the following experiments.
@@
The LPECs are solved via the MILP formulation in (18) and solve it with the commercial solver Gurobi Gurobi or the open-source solver HiGHS Huangfu2018 . In our experiments, the MILP reformulation was significantly faster that regularization- and penalty-based methods, cf. Appendix C. For the MILPs, we set the maximum number of nodes in the branch and bound algorithm to 5000 and the maximum computation time to 5 minutes. The tolerance for the LP solves in both MILP solvers is set to $10^{-9}$ (default is $10^{-6}$).
For comparisons, we also implement the regularization and penalty-based methods. Thereby, an adaptation of Reg($\tau$), $\ell_{1}$ and $\ell_{\infty}$ Kim2020 ; Ralph2004 ; Nurkanovic2024b to the form of  (38) are solved in a homotopy loop with a decreasing parameter $\tau$. We initialize $\tau^{0}=1$ and update it using the rule $\tau^{k}=0.1\tau^{k-1}$. The loop terminates if either: the NLP solution is feasible and the complementarity tolerance $h_{\perp}(x)$ is below $10^{-9}$; NLP is locally infeasible; or a maximum number of $15$ iterations is reached. A homotopy loop is declared successful if it finds a stationary point.
Additionally, we compare MPECopt to solving the MPEC directly as an NLP in 2 with IPOPT, and solving its MINLP reformulation (21) with a nonlinear branch-and-bound method implemented in Bonmin Bonami2008 .
We mention the remaining parameter choices in our implementation. The parameters in Algorithm 2 are $\kappa=0.1$, $\tau^{0}=1$, ${\rho}^{I}=10^{-1}$, and $\texttt{tol}_{h}=10^{-10}$. In the implementation of Algorithm 3, in line 6, a point is declared to be B-stationary if $\|d^{k,l}\|_{\infty}\leq\texttt{tol}_{\mathrm{B}}$. However, since there is also the trust region constraint $\|d\|\leq\rho^{k,l}$ we should avoid having $\rho^{k,l}<\texttt{tol}_{\mathrm{B}}$. Since the inner loop terminates within a finite number of iterations (cf. Proposition 4.3), this situation can be avoided by choosing appropriate parameters of the algorithm; we pick $\texttt{tol}_{\mathrm{B}}=10^{-8}$, $\rho^{k,0}=10^{-3}$, $N^{\mathrm{in}}=6$ and $\gamma^{\mathrm{L}}=0.1$. We also set a maximum number of outer iterations to $N^{\mathrm{out}}=25$. If no B-stationary point is found, the algorithm terminates after reaching the maximum number of iterations.
Our implementation of MPECopt is open-source and available at github.com/nosnoc/mpecopt. It contains code the reproduce the two benchmark results below, and several tutorial examples.
@@ 5.2 Results on MacMPEC benchmark
The MacMPEC problem set is available in the form of a package of AMPL .mod and .dat files from wiki.mcs.anl.gov/leyffer/index.php/MacMPEC. We consider in total 191 problems. Results of the experiments are given in the Dolan-Moré absolute and relative performance profiles Dolan2002 .
@@
Figure 5: Evaluating different MPEC solution methods on the MacMPEC test set in terms of finding a B-stationary point.
We compare MPECopt to other mentioned approaches. For Phase I of MPECopt, we use Algorithm (2) with Gurobi as LPEC solvers, as it had the highest success compared to other variants we tested. In the experiments, we compare in total six algorithms:
@@
    MPECopt, with $\ell_{\infty}$-penalty-based Phase I, and Gurobi as LPEC solver – MPECopt-$\ell_{\infty}$-Gurobi.
@@
    Regularization method that solve Reg($\tau$) in a homotopy loop – Reg.
@@
    $\ell_{\infty}$-penalty method, that solve an $\ell_{\infty}$ reformulation of (1Kim2020 in a homotopy loop – $\ell_{\infty}$.
@@
    Directly solving the MPEC as an NLP (2) with IPOPTNLP.
  6. 6.
    Solving the MINLP reformulation (21) with a nonlinear branch and bound in BonminMINLP.
@@
We also tried the $\ell_{1}$ penalty method, which similar computation time as $\ell_{\infty}$, but a slightly lower success rate. Moreover using HiGHS instead of Gurobi lead to slightly slower method and marginally lower success rate.
Therefore, for better readability, we omit these variants from the plots. Since all method use IPOPT as NLP solver, we can directly compare the computation times. Fig. 5 reports the absolute and relative performance profiles for this experiment.
@@
Note that, if MPECopt is successful, it always returns a B-stationary point, whereas for the other methods, we can only know that it is a B-stationary point if it is also S-stationary. To verify B-stationarity of the solutions computed by all other methods, we proceed as described in Appendix B.
We do not add the computational cost of verifying B-stationarity to the overall computation time of the other methods.
We observe that MPECopt-Reg-Gurobi achieves the highest overall success rate of 94.24%. It is both faster and more successful than its counterpart Reg (93.2%). The $\ell_{\infty}$-penalty method is the fastest overall and faster than MPECopt-Reg-Gurobi, but it attains a lower success rate of 92.14%. Meanwhile, MPECopt-$\ell_{\infty}$-Gurobi is only slightly slower than $\ell_{\infty}$ and achieves the same success rate. As expected, the plain NLP and MINLP reformulations are less competitive than the MPEC-tailored methods, with success rates of 90.00% and 72.52%, respectively. The main reasons for failure are timeouts and unsuccessful restoration phases. Additionally, the MINLP method frequently terminates at feasible but non–B-stationary points (39 cases).
The MPECopt method inherits the efficacy of the standard methods used in Phase I but can often be even faster, as it requires solving fewer NLPs. This is because, already for moderate values of $\tau$, the LPEC can identify a feasible BNLP in Phase I. Consequently, MPECopt needs to solve fewer NLPs than Reg and $\ell_{\infty}$. The branch-and-bound MINLP method solves the largest number of NLPs, which also explains its slow computation times.
In practice, only a few LPECs need to be solved. For instance, on the MacMPEC benchmark, the number of LPECssolved ranged from 1 to 6. In many cases, only two are required: one in Phase I to identify a feasible BNLP, and one in Phase II to verify B-stationarity.
We prove in Section 4 that, unless we need to verify B-stationarity, the LPECs do not need to be solved to global optimality; finding a feasible point is sufficient. Moreover, Proposition 3.1 shows that if the solution is S-stationary, it can be verified by solving a single LP. We examine these theoretical results by plotting histograms of the number of nodes in a MILP call (corresponding to the number of LPs solved) in both Phase I and Phase II. Fig. 6a shows the histogram when the MILP is solved to global optimality, while Fig. 6b shows the case when it is terminated early, at a feasible point. In both cases, Phase II only needed to solve LPs, and in Phase I we observe a significant reduction in the number of LPs to be solved in several cases. In fact, with early termination, on MacMPEC we only need to solve LPs. This illustrates our theoretical results: early termination of the MILP method reduces the number of LPs that must be solved. This did not significantly affect the overall computation time, as the largest computational burden on the MacMPEC problem set came from solving the NLP subproblems.
(a) Solving to optimality.
(b) Early termination.
Figure 6: Number of nodes in the MILP solver in Phase I and $I$. The node counter on the horizontal axis is on logarithmic scale.
@@
In summary, all methods – sometimes with minor tuning – can successfully compute B-stationary points on most MacMPEC problems, although only MPECopt directly certifies this. All methods are quite fast on easy problems without biactive constraints, where often a single NLP solve is sufficient. In fact, more than two-thirds of the solutions found by any algorithm correspond to problems with an empty biactive set. To provide a broader perspective, we present another benchmark with larger problems in the next section.